# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v2

@@ B.4 Modified log-linear learning in symmetric potential games
In the following, we consider a symmetric potential game and show that in this setting, we obtain convergence time guarantees to an $\epsilon$-efficient NE that depend polynomially on the number of players $N$ and on $1/\epsilon$. This improves the convergence time result provided in [32] which shows an exponential dependence on $1/\epsilon$.
@@
$\Psi^{\mathcal{A}}:=\bigg{\{}\bigg{(}\frac{v_{1}}{N},\ldots,\frac{v_{A}}{N}\bigg{)}\>|\>v_{j}\in\mathbb{Z}_{+}\>\forall j\in[A],\>\sum_{j=1}^{A}v_{j}=N\bigg{\}}$
@@
To obtain their results for symmetric potential games, Shah and Shin [32] propose modified log-linear learning, a variant of log-linear learning. In modified log-linear learning every player $i$ has an independent exponential clock of rate $\alpha/z_{i}^{t}$, where $\alpha>0$ is a parameter and $z_{i}^{t}:=1/N|\{j\in[N]\>|\>a_{j}^{t}=a_{i}^{t}\}|$. [^5] The only change compared to log-linear learning is that in log-linear learning players have a fixed exponential clock rate of $1$. This means that the times between two consecutive clock-ticks are independent and distributed as the exponential distribution of mean $\alpha/z_{i}^{t}$. When the clock of player $i$ ticks, she is allowed to alter her current action. Player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ defined as in Equation (3). Modified log-linear learning induces an aperiodic and irreducible Markov chain on the lower-dimensional state space $\Psi^{\mathcal{A}}$ with stationary distribution $\mu_{m}\in\Delta(\Psi^{\mathcal{A}})$ given by [32, Lemma 2]:
@@
$t\geq\frac{N}{\alpha c}e^{3\beta}\left(\log((A-1)\log(N+1))+\log\beta+2\log\frac{4}{\epsilon}\right)=\tilde{\Omega}\left(N\left(\frac{N^{A}}{\epsilon}\right)^{\frac{1}{\max\{\epsilon,\Delta\}}}\right),$
@@
Proof.
Define the set $\Psi_{*}^{\mathcal{A}}=\{x^{*}\in\Psi^{\mathcal{A}}\>|\>x^{*}\in\operatorname*{arg\,max}_{x\in\Psi^{\mathcal{A}}}\!\!\Phi(x)\}$ as the set of potential maximizers with cardinality $\!Y_{*}\!=\!|\Psi_{*}^{\mathcal{A}}|$ and define the set $\Psi^{\mathcal{A}}(\epsilon)=\{x\in\Psi^{\mathcal{A}}\>|\>\Phi(x)\geq\Phi(x^{*})\}$ as the set of $\epsilon$-approximate potential maximizers with cardinality $\!Y(\epsilon)\!=\!|\Psi^{\mathcal{A}}(\epsilon)|$. Then, the expected value of the potential function $\Phi(\cdot)$ over the stationary distribution $\mu_{m}$ of modified log-linear learning in (5) can be bounded as follows:
@@
The proof now follows from the same analysis as in the proof of Theorem $3$ in [32] with the exception that we do not make use of Lemma $6$ in [32] but replace it with our analysis above. Concretely, we set $\beta$ as specified in Equation (22) rather than as in [32, Eq. (8)]. ∎
@@ Appendix C Binary log-linear learning
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The additional term $1/A$ stems from the fact that player $i$ first randomly samples an action $\tilde{a}_{i}$ and then decides between this action and her previous action. Arslan et al. [1] show that its stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ is given by :
@@
The stationary distribution of binary log-linear learning is the same as that of log-linear learning (Equation (5)). Thus, log-linear- and binary log-linear learning converge to an approximately efficient Nash equilibrium in the long run. We briefly outline the proof of Theorem 2 and then provide a detailed proof.
Proof outline.
The proof follows from the same line of arguments as the proof of Theorem 1. In particular, the first step in the proof of Theorem 1 remains the same since binary log-linear learning has the same stationary distribution as log-linear learning. In the second step in the proof of Theorem 1, the main difference is that the transition matrix in (23) of binary log-linear learning differs from the transition matrix in (4) of log-linear learning. Thus, the log-Sobolev constant of binary log-linear can be lower-bounded as follows:
@@
Then, we use Lemma 1 to show that
@@
for $t\geq\frac{1}{\rho(PP^{*})}(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon})$ with $\rho(PP^{*})$ lower-bounded as in Equation (25). ∎
Proof of Theorem 2.
By Lemma 3, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
@@
Equation (2) in Lemma 1 provides the following upper bound on the mixing time:
@@
where the third line follows from Lemma 2, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$ for $t$ set as in Equation (26), and the fact that $\Phi(\cdot)\in[0,1]$. Lemma 2 is applicable when all players adhere to binary-based log-linear learning rather than log-linear learning since the proof of Lemma 2 depends only on the stationary distribution $\mu$ of the corresponding learning rule which is the same for log-linear learning and binary log-linear learning. This concludes the proof of Theorem 2. ∎
@@ Appendix D Robustness of log-linear learning
In this section, we prove Theorem 3 and apply it to the corrupted-utility case to prove Corollary 2.
D.1 Proof of Theorem 3
Proof of Theorem 3.
Consider a learning rule $P$, to prove Theorem 3, we first provide a decomposition that relates the expected value of the potential when the agents follow $P$ defined in Equation (8) to the same quantity when the agents follow $P_{\ell}$ defined in Equation (4) instead.
@@ D.1 Proof of Theorem 3
where the last line follows because $|\Phi(a)|\leq 1$ for all $a\in\mathcal{A}^{N}$ and because .$\|.\|_{1}\leq\sqrt{A^{N}}\|\cdot\|_{2}$.
The rest of the proof is based on controlling $\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ using our mixing time results.
Decomposition: Using Lemma 4 we obtain:
@@
where $L(P_{\ell})=\frac{2A^{N}}{\rho(P_{\ell}P_{\ell}^{*})}(\log\log\frac{1}{\mu_{\ell,\min}}+\log(8A^{N}))$ follows from Lemma 4. In Theorem 1 we showed that $\mu_{\ell,\min}\geq\frac{e^{-\beta}}{A^{N}}$ and $\rho(P_{\ell}P_{\ell}^{*})\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}}$, therefore
@@
$\begin{cases}t&\geq t_{\operatorname{mix}}^{P}(\epsilon/(2\sqrt{A^{N}})))\\ t^{\prime}&\to\infty\\ \beta&=\log\bigg{(}(A^{N}-A^{N}(\epsilon))\bigg{(}\frac{2}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\bigg{)}\bigg{)}/\max\{\epsilon,\Delta\}\end{cases}$
@@
$\begin{cases}\|P^{t}-P_{\ell}^{t}\|_{2}&\leq\epsilon/\left(2\sqrt{A^{N}}\right)+L(P_{\ell})\|P_{1}-P_{\ell}\|_{2}\\ \mathbb{E}_{\mu^{0}P_{\ell}^{t^{\prime}}}[\Phi]&\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2,\\ L(P_{\ell})&=\mathcal{O}\Bigg{(}N^{2}A^{N+5}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}\Bigg{(}\log\log A^{N}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}+\log(A^{N})\Bigg{)}\Bigg{)},\end{cases}$
where the second line follows from Lemma 2. Plugging the above inequalities with the bound on $L$ from Lemma 4 into the decomposition (28) proves the desired result.
We now bound the mixing time $t_{\operatorname{mix}}^{P}(\epsilon/(2\sqrt{A^{N}}))$ of the Markov chain induced by $P$.
Mixing time bound: To bound the mixing time of the Markov chain induced by $P$, we use inequality (2) and Lemma 3. Assuming a lower bound of $p_{\min}/N$ on the probabilities of all feasible transitions implies a bound on the stationary distribution as we will show next.
@@
Deducing the mixing-time bound: We can now give an explicit bound on the mixing time of the transition $P$. First, we have by Lemma 3:
@@
$t_{\operatorname{mix}}^{P}(\epsilon/(2\sqrt{A^{N}}))$ $\leq\frac{1}{\rho(PP^{*})}\Bigg{(}\log\log\frac{1}{(\mu_{P})_{\min}}+2\log\frac{2\sqrt{A^{N}}}{\epsilon}\Bigg{)}$
$\leq\frac{25N^{3/2}e^{N}}{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}\Bigg{(}\log\log\frac{e^{N}}{p_{\min}^{N}\sqrt{2\pi N}}+2\log\frac{2\sqrt{A^{N}}}{\epsilon}\Bigg{)}.$
This concludes the proof of Theorem 3. ∎
D.2 Proof of Corollary 2
@@
Also, since $P_{a,\tilde{a}}\geq P_{a,\tilde{a}}e^{-2\beta\xi}$ and using $P_{a,\tilde{a}}\geq\frac{e^{-\beta}}{NA}$ then we deduce that $P_{a,\tilde{a}}\geq\frac{e^{-\beta(1+2\xi)}}{NA}$. We conclude the proof with a straightforward application of Theorem 3 with $p_{\min}=e^{-\beta(1+2\xi)}/A$ and $\|\hat{P}-P\|_{2}\leq\frac{7A^{N}}{2N}\beta\xi$. ∎
D.3 Proof of Corollary 3
Similar to Corollary 2, we proceed by showing that the transition matrix of the Markov chain induced by fixed-share log-linear learning is close to that of log-linear learning.
Proof of Corollary 3.
@@ D.3 Proof of Corollary 3
which entails that $\hat{P}$ satisfies the condition of Theorem 3 with $p_{\min}\geq\frac{\xi}{A}+\frac{(1-\xi)e^{-\beta}}{A}$.
Additionally, we can show that:
@@
where $P$ is the transition matrix of log-linear learning.
Therefore,
@@
The proof is then concluded by a straightforward application of Theorem 3 with $p_{\min}\geq\frac{\xi}{A}+\frac{(1-\xi)e^{-\beta}}{A}$ and $\|\hat{P}-P\|_{2}\leq\xi\sqrt{\frac{A^{N}}{N}}$.
D.4 Proof of Lemma 4
Proof of Lemma 4.
@@ D.4 Proof of Lemma 4
The proof is then concluded by using the mixing-time bound from inequality (2). ∎

### v3

@@ B.4 Modified log-linear learning in symmetric potential games
In the following, we consider a symmetric potential game and show that in this setting, we obtain convergence time guarantees to an $\epsilon$-efficient NE that depend polynomially on the number of players $N$ and on $1/\epsilon$. This improves the convergence time result provided in (Shah and Shin, 2010) which shows an exponential dependence on $1/\epsilon$.
@@
$\Psi^{\mathcal{A}}:=\bigg{\{}\bigg{(}\frac{v_{1}}{N},\ldots,\frac{v_{A}}{N}\bigg{)}\>|\>v_{j}\in\mathbb{Z}_{+}\>\forall j\in[A],\>\sum_{j=1}^{A}v_{j}=N\bigg{\}}$ (22)
@@
To obtain their results for symmetric potential games, Shah and Shin (2010) propose modified log-linear learning, a variant of log-linear learning. In modified log-linear learning every player $i$ has an independent exponential clock of rate $\alpha/z_{i}^{t}$, where $\alpha>0$ is a parameter and $z_{i}^{t}:=1/N|\{j\in[N]\>|\>a_{j}^{t}=a_{i}^{t}\}|$. [^8] The only change compared to log-linear learning is that in log-linear learning players have a fixed exponential clock rate of $1$. This means that the times between two consecutive clock-ticks are independent and distributed as the exponential distribution of mean $\alpha/z_{i}^{t}$. When the clock of player $i$ ticks, she is allowed to alter her current action. Player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ defined as in Equation (3). Modified log-linear learning induces an aperiodic and irreducible Markov chain on the lower-dimensional state space $\Psi^{\mathcal{A}}$ with stationary distribution $\mu_{m}\in\Delta(\Psi^{\mathcal{A}})$ given by (Shah and Shin, 2010, Lemma 2):
@@
$t\geq\frac{N}{\alpha c}e^{3\beta}\left(\log((A-1)\log(N+1))+\log\beta+2\log\frac{4}{\epsilon}\right)=\tilde{\mathcal{O}}\left(N\left(\frac{N^{A}}{\epsilon}\right)^{\frac{1}{\max\{\epsilon,\Delta\}}}\right),$
@@
Proof B.8.
[…]
@@
The proof now follows from the same analysis as in the proof of Theorem $3$ in (Shah and Shin, 2010) with the exception that we do not make use of Lemma $6$ in (Shah and Shin, 2010) but replace it with our analysis above. Concretely, we set $\beta$ as specified in Equation (23) rather than as in (Shah and Shin, 2010, Eq. (8)).
@@ Appendix C Binary log-linear learning
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The additional term $1/A$ stems from the fact that player $i$ first randomly samples an action $\tilde{a}_{i}$ and then decides between this action and her previous action. Arslan et al. (2007) show that its stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ is given by :
@@
The stationary distribution of binary log-linear learning is the same as that of log-linear learning (Equation (5)). Thus, log-linear- and binary log-linear learning converge to an approximately efficient Nash equilibrium in the long run. We briefly outline the proof of Theorem 4.1 and then provide a detailed proof.
Proof C.1.
The proof follows from the same line of arguments as the proof of Theorem 3.1. In particular, the first step in the proof of Theorem 3.1 remains the same since binary log-linear learning has the same stationary distribution as log-linear learning. In the second step in the proof of Theorem 3.1, the main difference is that the transition matrix in (24) of binary log-linear learning differs from the transition matrix in (4) of log-linear learning. Thus, the log-Sobolev constant of binary log-linear can be lower-bounded as follows:
@@
Then, we use Lemma 2.2 to show that
@@
for $t\geq\frac{1}{\rho(PP^{*})}(\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon})$ with $\rho(PP^{*})$ lower-bounded as in Equation (26).
Proof C.2.
of Theorem 4.1 By Lemma 3.4, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
@@
Equation (2) in Lemma 2.2 provides the following upper bound on the mixing time:
@@
where the third line follows from Lemma 3.3, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/4$ for $t$ set as in Equation (27), and the fact that $\Phi(\cdot)\in[0,1]$. Lemma 3.3 is applicable when all players adhere to binary-based log-linear learning rather than log-linear learning since the proof of Lemma 3.3 depends only on the stationary distribution $\mu$ of the corresponding learning rule which is the same for log-linear learning and binary log-linear learning. This concludes the proof of Theorem 4.1.
@@ Appendix D Robustness of log-linear learning
In this section, we prove Theorem 4.3 and apply it to the corrupted-utility case to prove Corollary 4.4.
D.1 Proof of Theorem 4.3
Proof D.1.
Consider a learning rule $P$, to prove Theorem 4.3, we first provide a decomposition that relates the expected value of the potential when the agents follow $P$ defined in Equation (8) to the same quantity when the agents follow $P_{\ell}$ defined in Equation (4) instead.
@@
where the last line follows because $|\Phi(a)|\leq 1$ for all $a\in\mathcal{A}^{N}$ and because .$\|.\|_{1}\leq\sqrt{A^{N}}\|\cdot\|_{2}$. The rest of the proof is based on controlling $\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ using our mixing time results.
Decomposition: Using Lemma 4.2 we obtain:
@@
where $L(P_{\ell})=\frac{2A^{N}}{\rho(P_{\ell}P_{\ell}^{*})}(\log\log\frac{1}{\mu_{\ell,\min}}+\log(8A^{N}))$ follows from Lemma 4.2. In Theorem 3.1, we showed that $\mu_{\ell,\min}\geq\frac{e^{-\beta}}{A^{N}}$ and $\rho(P_{\ell}P_{\ell}^{*})\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}}$, therefore
@@
$\begin{cases}t&\geq t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}})))\\ t^{\prime}&\to\infty\\ \beta&=\log\bigg{(}(A^{N}-A^{N}(\epsilon/2))\bigg{(}\frac{4}{\epsilon/2A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\bigg{)}\bigg{)}/\max\{\epsilon/2,\Delta\}\end{cases}$
@@
$\begin{cases}\|P^{t}-P_{\ell}^{t}\|_{2}&\leq\epsilon/\left(2\sqrt{A^{N}}\right)+L(P_{\ell})\|P-P_{\ell}\|_{2}\\ \mathbb{E}_{a\sim\mu^{0}P_{\ell}^{t^{\prime}}}[\Phi(a)]&\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2,\\ L(P_{\ell})&=\mathcal{O}\Bigg{(}N^{2}A^{N+5}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}\Bigg{(}\log\log A^{N}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}+\log(A^{N})\Bigg{)}\Bigg{)},\end{cases}$
where the second line follows from Lemma 3.3. Plugging the above inequalities with the bound on $L$ from Lemma 4.2 into the decomposition (29) proves the desired result. We now bound the mixing time $t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}}))$ of the Markov chain induced by $P$.
Mixing time bound: To bound the mixing time of the Markov chain induced by $P$, we use inequality (2) and Lemma 3.4. Assuming a lower bound of $p_{\min}/N$ on the probabilities of all feasible transitions implies a bound on the stationary distribution as we will show next.
@@
Deducing the mixing-time bound: We can now give an explicit bound on the mixing time of the transition $P$. First, we have by Lemma 3.4:
@@
$t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}}))$ $\leq\frac{1}{\rho(PP^{*})}\Bigg{(}\log\log\frac{1}{(\mu_{P})_{\min}}+2\log\frac{4\sqrt{A^{N}}}{\epsilon}\Bigg{)}$
$\leq\frac{25N^{3/2}e^{N}}{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}\Bigg{(}\log\log\frac{e^{N}}{p_{\min}^{N}\sqrt{2\pi N}}+2\log\frac{4\sqrt{A^{N}}}{\epsilon}\Bigg{)}.$
This concludes the proof of Theorem 4.3.
D.2 Proof of Corollary 4.4
@@
Also, since $P_{a,\tilde{a}}\geq P_{a,\tilde{a}}e^{-2\beta\xi}$ and using $P_{a,\tilde{a}}\geq\frac{e^{-\beta}}{NA}$ then we deduce that $P_{a,\tilde{a}}\geq\frac{e^{-\beta(1+2\xi)}}{NA}$. We conclude the proof with a straightforward application of Theorem 4.3 with $p_{\min}=e^{-\beta(1+2\xi)}/A$ and $\|\hat{P}-P\|_{2}\leq\frac{7A^{N}}{2N}\beta\xi$.
D.3 Proof of Corollary 4.5
Similar to Corollary 4.4, we proceed by showing that the transition matrix of the Markov chain induced by fixed-share log-linear learning is close to that of log-linear learning.
Proof D.3.
@@
which entails that $\hat{P}$ satisfies the condition of Theorem 4.3 with $p_{\min}\geq\frac{\xi}{A}+\frac{(1-\xi)e^{-\beta}}{A}$. Additionally, we can show that:
@@
where $P$ is the transition matrix of log-linear learning. Therefore,
@@
The proof is then concluded by a straightforward application of Theorem 4.3 with $p_{\min}\geq\frac{\xi}{A}+\frac{(1-\xi)e^{-\beta}}{A}$ and $\|\hat{P}-P\|_{2}\leq\xi\sqrt{\frac{A^{N}}{N}}$.
D.4 Proof of Lemma 4.2
Proof D.4.
@@
The proof is then concluded by using the mixing-time bound from inequality (2).
Appendix E Numerical Illustration
We illustrate our convergence time results for log-linear learning for two-player identical interest games. In our setup, each player has the action set $\mathcal{A}=\{1,2,\ldots,10\}$, and the players have the same utility matrix denoted by $\{U(a_{1},a_{2})\}_{a_{1},a_{2}\in[10]}$. We simulate $10$ different utility matrices $U(\cdot,\cdot)\in[0,1]^{10\times 10}$ of the following form: we first fix a suboptimality gap $\Delta$, then we set $U(1,1)=1$, $U(10,10)=1-\Delta$, and the remaining entries of the utility matrix are sampled uniformly from the range $[0,1-\Delta)$.
Figure 1: Expected potential value when all players follow log-linear learning with $\beta$ set as in Equation (6). Left: performance for $\epsilon=0.05$ and different values of $\Delta$. Right: performance for $\Delta=0.1$ and different values $\epsilon$. Lines are averages over 10 randomly generated games, shaded areas represent one standard deviation, and the stars mark the first time the desired precision $1-\epsilon$ is reached.
In our first experiment, we fix $\epsilon$ as $0.05$ and vary the suboptimality gap $\Delta$ in $[0.15,0.10,0.075]$. We plot the expected potential value as a function of time when all players follow log-linear learning with temperature parameter $\beta$ set as in Equation (6). Figure 1 (left) demonstrates that the convergence time increases as the suboptimality gap $\Delta$ decreases. This shows that the suboptimality gap $\Delta$ quantifies the convergence time of log-linear learning well and therefore should be taken into account when bounding the convergence time, as done in our Theorem 3.1. In addition, we observe that this increase in convergence time is super-linear which aligns with the polynomial convergence time of $\tilde{\mathcal{O}}((\frac{1}{\epsilon})^{1/\Delta})$ that we proved in Theorem 3.1 when $\epsilon$ is smaller than $\Delta$.
In our second experiment, we fix $\Delta$ as $0.1$ and vary the precision level $\epsilon$ in $[0.1,0.05,0.025,0.02]$. We then plot the expected potential value as a function of time when all players follow log-linear learning with temperature parameter $\beta$ set as in Equation (6). Figure 1 (right) shows that the convergence time increases as the value of $\epsilon$ decreases. In other words, the convergence time of log-linear learning depends inversely on the desired suboptimality. This is also reflected in our convergence time bounds of order $\tilde{\mathcal{O}}((\frac{1}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ in Theorem 3.1.